# Certified Defense on the Fairness of Graph Neural Networks

### v2

@@ 3.3 Certified Fairness Defense over Node Attributes and Graph Structure
Figure 1: An example about how ELEGANT works in the input space.
We let $\bm{\Gamma}_{\bm{A}}=\gamma_{\bm{A}}(\bm{\omega}_{\bm{A}},\mathcal{V}_{\text{vul}})$ below for simplicity. To better illustrate how classifier $\tilde{g}_{\bm{A},\bm{X}}$ achieves certified fairness defense over both data modalities of an attributed network, we provide an exemplary case in Figure 1. Here we assume node $v_{i}\in\mathcal{V}_{\text{vul}}$. Considering the high dimensionality of node attributes and adjacency matrix, we only analyze two entries $\bm{X}_{i,j}$ and $\bm{A}_{i,j}$ and omit other entries after noise for simplicity. Here the superscript $(i,j)$ represents the $i$-th row and $j$-th column of a matrix. Under binary noise, entry $\bm{A}_{i,j}$ only has two possible values, i.e., $\bm{A}_{i,j}\oplus 0$ and $\bm{A}_{i,j}\oplus 1$. We denote the two cases as Case (1) and Case (2), respectively. We assume that the area where $g$ returns 1 in the span of the two input random entries of $g$ (i.e., $\bm{X}_{i,j}$ and $\bm{A}_{i,j}$ under random noise) is an ellipse (marked out with green), where the decision boundary is marked out with deep green. In Case (1), $\bm{X}_{i,j}$ under random noise follows a Gaussian distribution, whose probability density function is marked out as deep red. We assume that, in this case, the integral of the probability density function within the range of the ellipse (marked out with shallow red) is larger than 0.5. Correspondingly, according to Definition 2, $\tilde{g}_{\bm{X}}$ returns 1 in this case. In Case (2), we similarly mark out the probability density function and the area used for integral within the range of the ellipse. We assume that in this case, the integral is smaller than 0.5, and thus $\tilde{g}_{\bm{X}}$ returns 0. Note that to compute the output of $\tilde{g}_{\bm{A},\bm{X}}$, we need to identify the output of $\tilde{g}_{\bm{X}}$ with the largest probability. Notice that $\beta>0.5$, we have that $\bm{A}_{i,j}\oplus 0$ happens with a larger probability than $\bm{A}_{i,j}\oplus 1$. Therefore, $\tilde{g}_{\bm{A},\bm{X}}$ outputs 1 in this example. In other words, the bias level of predictions of $f_{\bm{\theta}^{*}}$ is satisfying (i.e., smaller than $\eta$) based on $\tilde{g}_{\bm{A},\bm{X}}$.
@@
$\epsilon_{\bm{A}}=\max\tilde{\epsilon_{\bm{A}}},\;\text{s.t.}\;\underline{P_{\tilde{g}_{\bm{X}}=1}}>0.5,\;\forall\;\|\bm{\Delta}_{\bm{A}}\|_{0}\leq\tilde{\epsilon_{\bm{A}}}$ (1)
following the formulation given in [34, 4, 47]. Here $\underline{P_{\tilde{g}_{\bm{X}}=1}}$ is the lower bound of $\mathrm{Pr}(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Delta}_{\bm{A}}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})=1)$.
@@
$\epsilon_{\bm{X}}=\min\bigl{\{}\tilde{\epsilon_{\bm{X}}}:\tilde{\epsilon_{\bm{X}}}=\zeta\left(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})\right)\text{ , where}\;\bm{\Gamma}_{\bm{A}}\in\bar{\mathcal{A}}\bigl{\}}.$ (2)
@@
Here, under structure attacks, we follow the rationale of most works achieving certification in discrete space [34, 4, 47]: the probability of $\tilde{g}_{\bm{X}}$ will always be 1 (i.e., certification succeeds) if $\underline{P_{\tilde{g}_{\bm{X}}=1}}>0.5$. Meanwhile, taking the smallest $\epsilon_{\bm{X}}$ over all possible $\tilde{\epsilon_{\bm{X}}}$ will ensure the certification successful in the space of node attributes, since perturbations over node attributes smaller than $\epsilon_{\bm{X}}$ will not change the output of $\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})$ given any $\bm{\Gamma}_{\bm{A}}\in\bar{\mathcal{A}}$. We provide detailed proofs in Appendix A.
3.4 Certification in Practice
@@
Estimating the Predicted Label Probabilities. According to Definition 3, it is necessary to obtain $\mathrm{Pr}(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})=c)$ ($c\in\{0,1\}$) to determine the output of classifier $\tilde{g}_{\bm{X}}$. We propose to leverage a Monte Carlo method to estimate such a probability. Specifically, we first randomly pick $N$ samples of $\bm{\Gamma}_{\bm{A}}$ as $\bar{\mathcal{A}}^{\prime}$ ($\bar{\mathcal{A}}^{\prime}\subset\bar{\mathcal{A}}$). Considering the output of $\tilde{g}_{\bm{X}}$ is binary, we then follow a common strategy [8] to consider this problem as a parameter estimation of a Binomial distribution: we first count the number of returned label 1 and 0 under noise as $N_{1}$ and $N_{0}$ ($N_{1}+N_{0}=N$); then we choose a confidence level $1-\alpha$ and take the $\alpha$-th quantile of the beta distribution with parameters $N_{1}$ and $N_{0}$ as the estimated probability lower bound for returning label $c=1$. We proved that all theoretical analysis still holds true for such an estimation in Appendix. We follow a similar strategy to estimate the probability lower bound of yielding 1 for $g(\bm{A},\bm{X}+\bm{\Gamma}_{\bm{X}})$.
Obtaining Fair Classification Results. After achieving certified fairness defense based on $\tilde{g}_{\bm{A},\bm{X}}$, we also need to obtain the corresponding node classification results (given by $f_{\bm{\theta}^{*}}$) over $\mathcal{V}_{\text{tst}}$. We propose to collect all classification results associated with the sampled $\bm{\Gamma}_{\bm{A}}^{\prime}\in\bar{\mathcal{A}}^{\prime}$ that leads to an estimated lower bound of $\mathrm{Pr}(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}}^{\prime},\bm{X})=1)$ to be larger than 0.5 as $\hat{\mathcal{Y}}^{\prime}$. Here $\hat{\mathcal{Y}}^{\prime}$ is a set of output matrices of $f_{\bm{\theta}^{*}}$, where each matrix consists of the one-hot output classification results (as each row in the matrix) for all nodes. We propose to take $\mathrm{argmin}_{\hat{\bm{Y}}^{\prime}}\pi(\hat{\bm{Y}}^{\prime},\mathcal{V}_{\text{tst}}),s.t.\;\hat{\bm{Y}}^{\prime}\in\hat{\mathcal{Y}}^{\prime}$ as the final node classification results. Correspondingly, consider $\mathrm{Pr}(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}}^{\prime},\bm{X})=1)$ falls into the confidence interval characterized by $1-\alpha$, we have a neat probabilistic theoretical guarantee below.
Proposition 1.
Calculation of Perturbation Budgets. We calculate $\epsilon_{\bm{A}}$ with the algorithm proposed in [47]. For $\epsilon_{\bm{X}}$, we utilize a Monte Carlo method to estimate its value. Specifically, we leverage $\text{min}\{\tilde{\epsilon_{\bm{X}}}:\tilde{\epsilon_{\bm{X}}}\;\text{is derived with classifier}\;\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}}^{\prime},\bm{X})\text{ , where}\;\bm{\Gamma}_{\bm{A}}^{\prime}\in\bar{\mathcal{A}}^{\prime}\}$ to estimate the value of $\epsilon_{\bm{X}}$.
4 Experimental Evaluations
In this section, we aim to answer three research questions: RQ1: How well does ELEGANT perform in achieving certified fairness defense? RQ2: How does ELEGANT perform under fairness attacks compared to other popular fairness-aware GNNs? RQ3: How does ELEGANT perform under different settings of parameters? We present the main experimental settings and representative results in this section due to space limits. Detailed settings and supplementary experiments are in Appendix.
4.1 Experimental Settings
Downstream Task and Datasets. We focus on the widely studied node classification task, which is one of the most representative tasks in the domain of learning on graphs. We adopt three real-world network datasets that are widely used to perform studies on the fairness of GNNs, namely German Credit [1, 2], Recidivism [1, 26], and Credit Defaulter [1, 50]. We provide their basic information, including how these datasets are built and their statistics, in Appendix.
Evaluation Metrics. We perform evaluation from three main perspectives, including model utility, fairness, and certified defense. To evaluate utility, we adopt the node classification accuracy. To evaluate fairness, we adopt the widely used metrics $\Delta_{\text{SP}}$ (measuring bias under Statistical Parity) and $\Delta_{\text{EO}}$ (measuring bias under Equal Opportunity). To evaluate certified defense, we extend a traditional metric named Certified Accuracy [47, 8] in our experiments, and we name it as Fairness Certification Rate (FCR). Specifically, existing GNN certification works mainly focus on a certain individual node, and utilize certified accuracy to measure the ratio of nodes that are correctly classified and also successfully certified out of all test nodes [47]. In this paper, however, we perform certified (fairness) defense for individuals over an entire test set (instead of for any specific individual). Accordingly, we propose to sample multiple test sets out of nodes that are not involved in the training and validation set. Then we perform certified fairness defense for all sampled test sets, and utilize the ratio of test sets that are successfully certified over all sampled sets as the metric of certified defense. The rationale of FCR is leveraging a Monte Carlo method to estimate the probability of being successfully certified for a randomly sampled test node set.
GNN Backbones and Baselines. Note that ELEGANT serves as a plug-and-play framework for any optimized GNNs ready to be deployed. To evaluate the generality of ELEGANT across GNNs, we adopt three of the most representative GNNs spanning across simple and complex ones, namely Graph Sample and Aggregate Networks [18] (GraphSAGE), Graph Convolutional Networks [32] (GCN), and Jumping Knowledge Networks (JK). Note that to the best of our knowledge, existing works on fairness certification cannot certify the attacks over two data modalities (i.e., continuous node attributes and binary graph topology) at the same time, and thus cannot be naively generalized onto GNNs. Hence we compare the usability of GNNs before and after certification with ELEGANT. Moreover, we also adopt two popular fairness-aware GNNs as baselines to evaluate bias mitigation, including FairGNN [9] and NIFTY [1]. Specifically, FairGNN utilizes adversarial learning to debias node embeddings, while NIFTY designs regularizations to debias node embeddings.
Threat Models. We propose to evaluate the performance of ELEGANT and other fairness-aware GNN models under actual attacks on fairness. We first introduce the threat model over graph structure. To the best of our knowledge, FA-GNN [20] is the only work that performs graph structure attacks targeting the fairness of GNNs. Hence we adopt FA-GNN to attack graph structure. In terms of node attributes, to the best of our knowledge, no existing work has made any explorations. Hence we directly utilize gradient ascend to perform attacks. Specifically, after structure attacks have been performed, we identify the top-ranked node attribute elements (out of the node attribute matrix) that positively influence the exhibited bias the most via gradient ascend. For any given budget (of attacks) on node attributes, we add perturbations to these elements in proportion to their gradients.
Table 1: Comparison between vanilla GNNs and certified GNNs under ELEGANT over three popular GNNs across three real-world datasets. Here ACC denotes node classification accuracy, and E- prefix marks out the GNNs under ELEGANT with certification. $\uparrow$ denotes the larger, the better; $\downarrow$ represents the opposite. Numerical values are in percentage, and the best ones are in bold.
German Credit Recidivism Credit Defaulter
4.2 RQ1: Fairness Certification Effectiveness
To answer RQ1, we investigate the performance of different GNNs after certification across different real-world attributed network datasets over FCR, utility, and fairness. We present the experimental results across three GNN backbones and three real-world attributed network datasets in Table 1. Here bias is measured with $\Delta_{\text{SP}}$, and we have similar observations on $\Delta_{\text{EO}}$. We summarize the main observations as follows: (1) Fairness Certification Rate (FCR). We observe that ELEGANT realizes values of FCR around or even higher than 90% for all three GNN backbones and three attributed network datasets, especially for the German Credit dataset, where vanilla GNNs tend to exhibit a high level of bias. The corresponding intuition is that, for nodes in any randomly sampled test set, we have a probability around or higher than 90% to successfully certify the fairness level of the predictions yielded by the GNN model with our proposed framework ELEGANT. Hence ELEGANT achieves a satisfying fairness certification rate across all adopted GNN backbones and datasets. (2) Utility. We found that compared with those vanilla GNN backbones, certified GNNs with ELEGANT also exhibit comparable and even higher node classification accuracy values in all cases. Hence we conclude that our proposed framework ELEGANT does not significantly jeopardize the utility of the vanilla GNN models, and those certified GNNs with ELEGANT still bear a high level of usability in terms of node classification accuracy. (3) Fairness. Although the goal of ELEGANT is not debiasing GNNs, we observe that certified GNNs with ELEGANT achieve better performances in all cases in terms of algorithmic fairness compared with those vanilla GNNs. This demonstrates that the proposed framework ELEGANT also contributes to bias mitigation. We conjecture that such an advantage of debiasing could be a mixed result of (1) adding random noise on node attributes and graph topology (as in Section 3.2 and Section 3.3) and (2) the proposed strategy of obtaining fair classification results (as in Section 3.4). We provide a more detailed analysis in Section B.8.
Figure 2: The bias levels of GCN, E-GCN, FairGNN, and NIFTY under fairness attacks on German Credit. The shaded bar indicates that certified budget $\epsilon_{\bm{A}}\leq\|\bm{\Delta}_{\bm{A}}\|_{0}$ or $\epsilon_{\bm{X}}\leq\|\bm{\Delta}_{\bm{X}}\|_{2}$. The y-axis is in logarithmic scale for better visualization purposes.
4.3 RQ2: Fairness Certification under Attacks
To answer RQ2, we perform attacks on the fairness of GCN, E-GCN, FairGNN (with a GCN backbone), and NIFTY (with a GCN backbone). Considering the large size of the quadratic space spanned by the sizes of perturbations $\bm{\Delta}_{\bm{A}}$ and $\bm{\Delta}_{\bm{X}}$, we present the evaluation under four representative ($\|\bm{\Delta}_{\bm{A}}\|_{0}$, $\|\bm{\Delta}_{\bm{X}}\|_{2}$) pairs. We set the threshold for bias $\eta$ to be 50% higher than the fairness level of the vanilla GCN model on clean data, since it empirically helps to achieve a high certification success rate under large perturbations.
We present the fairness levels of the four models in terms of $\Delta_{\text{EO}}$ in Figure 2. Note that we utilize a vanilla GCN to predict the labels for test nodes to obtain the classification results discussed in Section 3.4, and we also have similar observations on other GNNs/datasets. (1) Fairness. We found that the GCN model with the proposed framework ELEGANT achieves the lowest level of bias in all cases of fairness attacks. This observation is consistent with the superiority in fairness found in Table 1, which demonstrates that the fairness superiority of ELEGANT maintains even under attacks within a wide range of attacking perturbation sizes. (2) Certification on Fairness. We now compare the performance of E-GCN across different attacking perturbation sizes. We observed that under relatively small attacking perturbation sizes, i.e., ($2^{0},10^{-1}$), ($2^{1},10^{0}$), and ($2^{2},10^{1}$), ELEGANT successfully achieves certification over fairness, and the bias level increases slowly as the size of attacks increases. Under relatively large attacking perturbation size, i.e., ($2^{3},10^{2}$), although the attacking budgets go beyond the certified budgets, GCN under ELEGANT still exhibits a fairness level far lower than the given bias threshold $\eta$, and the fairness superiority maintains. Hence the adopted estimation strategies are safe in achieving fairness certification.
(a) FCR of certification for $\sigma$ over node attributes
(b) FCR of certification for $\beta$ over graph topology
Figure 3: Parameter study of $\sigma$ over $\epsilon_{\bm{X}}$ (a) and $\beta$ over $\epsilon_{\bm{A}}$ (b). Experimental results are presented based on GCN over German credit and Credit Defaulter for (a) and (b), respectively. Similar tendencies can also be observed based on other GNNs and datasets.
4.4 RQ3: Parameter Study
To answer RQ3, we propose to perform parameter study focusing on two most critical parameters, $\sigma$ and $\beta$. To examine how $\sigma$ and $\beta$ influence the effectiveness of ELEGANT in terms of both FCR and certified defense budgets, we set numerical ranges for $\epsilon_{\bm{X}}$ (from 0 to 1e1) and $\epsilon_{\bm{A}}$ (from 0 to $2^{6}$) and divide the two ranges into grids. In both ranges, we consider the dividing values of the grids as thresholds for certification budgets. In other words, under each threshold, we only consider the test sets with the corresponding certified defense budget being larger than this threshold as successfully certified ones, and the values of FCR are re-computed accordingly. Our rationale here is that with the thresholds (for $\epsilon_{\bm{X}}$ and $\epsilon_{\bm{A}}$) increasing, if FCR reduces slowly, this demonstrates that most successfully certified test sets are associated with large certified defense budgets. However, if FCR reduces fast, then most successfully certified test sets only bear small certified defense budgets.
Here we present the experimental results of $\sigma$ and $\beta$ with the most widely used GCN model based on German Credit in Figure 3(a) and Credit Defaulter in Figure 3(b), respectively. We also have similar observations on other GNNs and datasets. We summarize the main observations as follows: (1) Analysis on $\sigma$. We observe that most cases with larger $\sigma$ are associated with a larger FCR compared with the cases where $\sigma$ is relatively small. In other words, larger values of $\sigma$ typically make FCR reduce slower w.r.t. the increasing of $\epsilon_{\bm{X}}$ threshold. This indicates that increasing the value of $\sigma$ helps realize larger certified defense budgets on node attributes, i.e., the increase of $\sigma$ dominates the tendency of $\epsilon_{\bm{X}}$ given in Equation 2. Nevertheless, it is worth mentioning that if $\sigma$ is too large, the information encoded in the node attributes could be swamped by the Gaussian noise and finally corrupt the classification accuracy. Hence moderately large values for $\sigma$, e.g., 5e-1 and 5e0, are recommended. (2) Analysis on $\beta$. We found that (1) for cases with relatively large $\beta$ (e.g., 0.8 and 0.9), the FCR also tends to be larger (compared with cases where $\beta$ is smaller) at $\epsilon_{\bm{A}}$ threshold being 0. Such a tendency is reasonable, since in these cases, the expected magnitude of the added Bernoulli noise is small. Correspondingly, GNNs under ELEGANT perform similarly to vanilla GNNs, and thus an $\eta$ larger than the bias level of vanilla GNNs is easier to be satisfied (compared with cases under smaller values of $\beta$); (2) for cases with relatively large $\beta$, the value of FCR generally reduces faster (w.r.t. $\epsilon_{\bm{A}}$ threshold) than cases where $\beta$ is smaller. Therefore, we recommend that for any test set of nodes: (1) if the primary goal is to achieve certification with a high probability, then larger values for $\beta$ (e.g., 0.8 and 0.9) would be preferred; (2) if the goal is to achieve certification with larger budgets on the graph topology, smaller values for $\beta$ (e.g., 0.6 and 0.7) should be selected.
5 Related Work
Algorithmic Fairness in GNNs. Existing GNN works on fairness mainly focus on group fairness and individual fairness [12]. Specifically, group fairness requires that each demographic subgroup (divided by sensitive attributes such as gender and race) in the graph should have their fair share of interest based on predictions [36]. Adversarial training is among the most popular strategies [9, 12]. In addition, regularization [1, 15, 51], topology modification [11, 45], and orthogonal projection [40] are also commonly used strategies. On the other hand, individual fairness it requires that similar individuals should be treated similarly [13], where such similarity may be determined in different ways [27, 10]. Designing optimization regularization terms to promote individual fairness for GNNs is a common strategy [15, 10, 44]. Nevertheless, despite the research advancements in the field of algorithmic fairness on GNNs, the adversarial defense against fairness attacks still remains in its infancy and has not been thoroughly explored. To the best of our knowledge, our paper serves as the first comprehensive study dedicated to addressing this important research problem, paving the way for future investigations in this under-explored area.
Certified Defense in Discrete Input Spaces. Recent studies have proposed certified defenses that guarantee the same predicted label on clean and perturbed input data under limited perturbation budgets. One prominent approach is randomized smoothing [8], which enhances the robustness of classifiers by introducing Gaussian noise. However, in discrete input spaces such as graph structures, Gaussian noise can lead to loose certification bounds [34]. To address this issue, a pioneering method achieving certified defense in discrete input spaces (i.e., point-wise certification) [34] was proposed for binary classification tasks, aiming to tighten the certification bound in discrete spaces. It partitions the discrete input space into chunks with identical noise ratios. By accumulating the output probabilities across these chunks, a tight lower bound of the perturbed output probability can be achieved. If this lower bound exceeds $0.5$, the certification region can be properly characterized. Building upon this, subsequent studies have extended point-wise certification to different problem settings. One approach [4] focused on binary graph structures and introduced a data-dependent Bernoulli noise, which reduces the number of partitioned chunks. Given the sparsity of graph structures, this method improves the efficiency of point-wise certification for GNNs [34]. Another extension [47] adapted point-wise certification to multi-class classification. Different from these methods, our structure certification extends the approach in [4] with data-independent noise. In this way, such noise can be generated before the graph data is given, and thus it facilitates achieving a composition of certification for both binary graph structure and continuous node attributes.
6 Conclusion
In this paper, we study a novel problem of certifying GNN node classifiers on fairness. To address this problem, we propose ELEGANT, a framework designed to achieve certification on top of any optimized GNN node classifier associated with certain perturbation budgets, ensuring that it is impossible for attackers to degrade the fairness level of predictions within such budgets. Notably, ELEGANT is designed to serve as a plug-and-play framework for any optimized GNNs and does not rely on any assumptions regarding GNN structures or re-training processes. Extensive experiments verify the strong effectiveness and generalizability of ELEGANT across multiple GNN architectures and real-world datasets. While this paper primarily focuses on the widely studied node classification task, we also highlight the potential for extending this study to other graph-related tasks as a future research direction. We expect positive broader impacts including deploying fairness-safe GNNs in applications, and no significant negative broader impact needs to be highlighted here.
Appendix

### v3

@@ 3.3. Certified Defense over Node Attributes and Graph Structure
Figure 1. An example of how ELEGANT functions in the input space spanned by node attributes and graph structure.
We let $\bm{\Gamma}_{\bm{A}}=\gamma_{\bm{A}}(\bm{\omega}_{\bm{A}},\mathcal{V}_{\text{vul}})$ below for simplicity. To better illustrate how classifier $\tilde{g}_{\bm{A},\bm{X}}$ achieves certified fairness defense over both data modalities of an attributed network, we provide an exemplary case in Figure˜1. Here we assume node $v_{i}\in\mathcal{V}_{\text{vul}}$. Considering the high dimensionality of node attributes and adjacency matrix, we only analyze two entries $\bm{X}_{i,j}$ and $\bm{A}_{i,j}$ and omit other entries after noise for simplicity. Here the superscript $(i,j)$ represents the $i$-th row and $j$-th column of a matrix. Under binary noise, entry $\bm{A}_{i,j}$ only has two possible values, i.e., $\bm{A}_{i,j}\oplus 0$ and $\bm{A}_{i,j}\oplus 1$. We denote the two cases as Case (1) and Case (2), respectively. We assume that the area where $g$ returns 1 in the span of the two input random entries of $g$ (i.e., $\bm{X}_{i,j}$ and $\bm{A}_{i,j}$ under random noise) is an ellipse (marked out with green), where the decision boundary is marked out with deep green. In Case (1), $\bm{X}_{i,j}$ under random noise follows a Gaussian distribution, whose probability density function is marked out as deep red. We assume that, in this case, the integral of the probability density function within the range of the ellipse (marked out with shallow red) is larger than 0.5. Correspondingly, according to Definition˜2, $\tilde{g}_{\bm{X}}$ returns 1 in this case. In Case (2), we similarly mark out the probability density function and the area used for integral within the range of the ellipse. We assume that in this case, the integral is smaller than 0.5, and thus $\tilde{g}_{\bm{X}}$ returns 0. Note that to compute the output of $\tilde{g}_{\bm{A},\bm{X}}$, we need to identify the output of $\tilde{g}_{\bm{X}}$ with the largest probability. Notice that $\beta>0.5$, we have that $\bm{A}_{i,j}\oplus 0$ happens with a larger probability than $\bm{A}_{i,j}\oplus 1$. Therefore, $\tilde{g}_{\bm{A},\bm{X}}$ outputs 1 in this example. In other words, the bias level of predictions of $f_{\bm{\theta}^{*}}$ is satisfying (i.e., smaller than $\eta$) according to the definition of $\tilde{g}_{\bm{A},\bm{X}}$.
@@
(1) $\epsilon_{\bm{A}}=\max\tilde{\epsilon_{\bm{A}}},\;\text{s.t.}\;\underline{P_{\tilde{g}_{\bm{X}}=1}}>0.5,\;\forall\;\|\bm{\Delta}_{\bm{A}}\|_{0}\leq\tilde{\epsilon_{\bm{A}}}$
following the formulation given in (lee2019tight; bojchevski2020efficient; wang2021certified). Here $\underline{P_{\tilde{g}_{\bm{X}}=1}}$ is the lower bound of $\mathrm{Pr}(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Delta}_{\bm{A}}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})=1)$.
@@
(2) $\epsilon_{\bm{X}}=\min\bigl\{\tilde{\epsilon_{\bm{X}}}:\tilde{\epsilon_{\bm{X}}}=\zeta\left(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})\right)\text{ , where}\;\bm{\Gamma}_{\bm{A}}\in\bar{\mathcal{A}}\bigl\}.$
@@
Here, under structure attacks, we follow the rationale of most works achieving certification in discrete space (lee2019tight; bojchevski2020efficient; wang2021certified): the probability of $\tilde{g}_{\bm{X}}$ will always be 1 (i.e., certification succeeds) if $\underline{P_{\tilde{g}_{\bm{X}}=1}}>0.5$. Meanwhile, taking the smallest $\epsilon_{\bm{X}}$ over all possible $\tilde{\epsilon_{\bm{X}}}$ will ensure the certification successful in the space of node attributes, since perturbations over node attributes smaller than $\epsilon_{\bm{X}}$ will not change the output of $\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})$ given any $\bm{\Gamma}_{\bm{A}}\in\bar{\mathcal{A}}$. We provide detailed proofs in Appendix in the online version.
3.4. Certification in Practice
We now introduce how to bridge the gaps between theoretical design and practical feasibility. We show the algorithmic routine of the proposed ELEGANT in Algorithm 1.
@@ 3.4. Certification in Practice
Estimating the Predicted Label Probabilities. According to Definition˜4, it is necessary to obtain $\mathrm{Pr}(\tilde{g}_{\bm{X}}(\bm{A}\oplus\bm{\Gamma}_{\bm{A}},\bm{X})=c)$ ($c\in\{0,1\}$) to determine the output of classifier $\tilde{g}_{\bm{X}}$. We propose to leverage a Monte Carlo method to estimate such a probability. Specifically, we first randomly pick $N$ samples of $\bm{\Gamma}_{\bm{A}}$ as $\bar{\mathcal{A}}^{\prime}$ ($\bar{\mathcal{A}}^{\prime}\subset\bar{\mathcal{A}}$). Considering the output of $\tilde{g}_{\bm{X}}$ is binary, we then follow a common strategy (cohen2019certified) to consider this problem as a parameter estimation of a Binomial distribution: we first count the number of returned label 1 and 0 under noise as $N_{1}$ and $N_{0}$ ($N_{1}+N_{0}=N$); then we choose a confidence level $1-\alpha$ and take the $\alpha$-th quantile of the beta distribution with parameters $N_{1}$ and $N_{0}$ as the estimated probability lower bound for returning label $c=1$. We proved that all analysis remains true for such an estimation in Appendix. Similar strategy is used to estimate the probability lower bound of yielding 1 for $g(\bm{A},\bm{X}+\bm{\Gamma}_{\bm{X}})$.
Algorithm 1 Certified Defense on the Fairness of GNNs